OpenStudy (anonymous):

how to find the common ratio of a geometric sequence if a1=40, n=4, and an=5 thank you all!

OpenStudy (anonymous):

Formula: a1x r^n-1= an This simplifies to 40 x r^4-1 = 5 40r^3=5 r = ?

OpenStudy (anonymous):

r=1/2! thank you so much!